# TrackID3x3: A Dataset and Algorithm for Multi-Player Tracking with Identification and Pose Estimation in 3x3 Basketball Full-court Videos

### v1

@@ 4.4 Pose Estimation Baselines
In all datasets, we evaluate the performance of RTMPose [35], HRNet [73], and SwinPose [87] to estimate the human pose. These models are common baseline models for top-down 2D pose estimation with diverse architectures.
@@
We use The Percentage of Detected Joints (PDJ) [79], widely used for evaluating the accuracy of 2D human pose estimation. It measures the proportion of correctly detected keypoints based on a distance threshold relative to the torso diameter. This metric is useful for assessing pose estimation models with varying degrees of localization precision [90]. In this evaluation metric, a keypoint is deemed correctly detected if the normalized distance between the predicted and ground truth keypoints is less than 0.5. Furthermore, PDJ allows for the evaluation of pose estimation accuracy under different levels of localization precision by adjusting the threshold. The area under the PDJ curve (AUC) is computed to assess performance across various threshold values.
@@ 5 Experiments
5.1 Performance in Track-ID task
Experiments were conducted on the Track-ID task on fixed camera datasets, Indoor and Outdoor, as described above. We experimented with this task mainly on the in-play portion: Indoor was all frames of each video, while Outdoor excluded the out-of-play portion and divided the video of one game into several segments. Two types of video segments exist: one starting with a check-ball at the top and the other starting with a free throw. The numbers of video frames in Indoor and Outdoor datasets 179.38 $\pm$ 64.77 and 1202.84 $\pm$ 1025.92 (mean $\pm$ standard deviation), respectively. The evaluation metrics are TI-HOTA and its sub-scores DetA, AssA, FN, and FP as described above. For the distance tolerance parameter, $\tau$ of TI-HOTA, $\tau=0.5$ was adopted given the ratio to the football pitch.
Table 3 shows the results with means and standard deviations; it is difficult to simply compare the Indoor and Outdoor results, but the better performance of Indoor can be explained to some extent by the shorter video length and the easier identification of attributes. This is because if ID exchange between players occurs in tracking, the attributes information will no longer be correct from then on, and the match between the prediction and the ground truth will fail for a large part of the video (i.e. FP, FN will increase and DetA will decrease).
@@ 5.1 Performance in Track-ID task
Indoor RTMPose [35] $39.82$% $87.26$% $71.99$% $70.02$% $77.87$% $78.29$% $73.28$% $36.90$%
HRNet [73] $40.71$% $86.76$% $68.62$% $66.36$% $77.31$% $75.60$% $77.26$% $36.06$%
SwinPose[87] $40.33$% $86.13$% $64.19$% $67.48$% $77.27$% $74.44$% $71.21$% $35.86$%
Outdoor RTMPose [35] $88.82$% $93.57$% $91.03$% $85.86$% $89.40$% $87.60$% $89.43$% $45.12$%
HRNet [73] $88.39$% $93.88$% $89.53$% $85.65$% $91.43$% $89.09$% $88.51$% $45.21$%
SwinPose[87] $88.21$% $93.14$% $89.19$% $85.20$% $90.16$% $90.16$% $89.27$% $45.03$%
Drone RTMPose [35] $85.57$% $85.60$% $85.17$% $80.64$% $77.59$% $84.30$% $83.45$% $42.07$%
HRNet [73] $85.99$% $86.09$% $84.53$% $79.69$% $84.60$% $84.75$% $84.07$% $42.38$%
SwinPose[87] $85.49$% $86.30$% $84.43$% $79.72$% $83.05$% $84.70$% $83.88$% $42.28$%
@@
Table 5: 2D pose estimation models performance on all datasets, ranked by PDJ. The best-performing result for each dataset is highlighted in bold. Sho, El, Wri, Cen, and Ank represent the average PDJ of Shoulder, Elbow, Wrist, Center, and Ankle, respectively. The Center is defined as the midpoint between the left and right hips. The PDJ column represents the mean PDJ across all 10 keypoints. AUC represents the area under the PDJ curve as the threshold varies from 0 to 0.5 in increments of 0.01, with a maximum value of 0.5.
5.2 Tracking Performance in Drone Dataset
Table 4 shows the results using four sequences of the Drone dataset with the default parameters of the two methods. As the frames contained numerous non-athlete individuals, including referees and spectators, we subsequently performed manual filtering to remove irrelevant bboxes, retaining only those of the six basketball players. The results reveal that the BoT-SORT method demonstrates notable improvement in HOTA scores compared to ByteTrack, along with a reduction in ID switches. However, there remains considerable room for further performance enhancement due to many occlusions.
5.3 2D Pose Estimation Performance
@@ 5.2 Tracking Performance in Drone Dataset
We also evaluated the performance of 2D pose estimation models in our all datasets. We evaluated three models without fine-tuning as baselines. Since each player in our datasets is annotated with only 10 keypoints, while the baseline models output 17 keypoints in COCO [45] keypoints format, we select the corresponding 10 keypoints for evaluation to ensure consistency.
@@
The evaluation results are summarized in Table 5. All models achieved their highest PDJ scores on the Outdoor dataset, but performance declined on the Drone dataset, with the Indoor dataset exhibiting the lowest accuracy. The lower performance observed in the Indoor dataset can be attributed to two primary factors. First, the image resolution of the Indoor dataset is 1280 × 720, which is significantly lower than that of the Outdoor and Drone datasets (3820 × 2160). The reduced resolution may lead to blurred keypoints and reduced model accuracy. Second, the PDJ scores for head keypoints in the Indoor dataset are considerably lower. This discrepancy arises because the head keypoints in Indoor data are annotated on the upper forehead rather than the nose, which introduces a larger localization error than other keypoints.
@@
A comparison between different keypoints across datasets, ankles, shoulders, and the center exhibit the most stable and high detection accuracy in all scenarios. In contrast, the elbows and wrists in the upper limbs demonstrate weaker performance, probably due to occlusion and motion blur. In our TrackID3x3 dataset, RTMPose[35] achieved the best performance in the Outdoor dataset, with an average PDJ of 89.43%. This result is highly comparable to the average PDJ of 89.51% reported for RTMPose in AutoSoccerPose [90]. For illustration, Figure 3 presents examples of players with lower prediction accuracy in the three datasets.
@@
Figure 3: Examples of players with low PDJ across three datasets. From left to right, the images correspond to the Indoor, Outdoor, and Drone datasets. The players positioned at the back are the primary targets for keypoint detection. However, due to occlusion, keypoint prediction for the obstructed player results in partial or complete failure.
6 Conclusion
@@
As a point to note in the experimental results, the biggest challenge in the Track-ID tasks and multi-object tracking in our dataset is ID switches due to frequent occlusions, especially track ID exchange between players or between players and referees, which is a major obstacle to accurate players’ data extraction. Addressing this challenge will be the highest priority. In addition, for the Track-ID task, both for prediction and ground truth, bbox-based position coordinates are not robust to arm extension and contraction and leg movement. This can be a hindrance in capturing the precise position and distance moved by the players. In future work, it may be a good idea to consider position coordinates that also utilize posture information [93].
Acknowledgements. This work was financially supported by JST PRESTO Grant Number JPMJPR20CA, JSPS KAKENHI Grant Number 23H03282, and Mizuno Sports Promotion Foundation.

### v2

@@ 4.3. Pose Estimation Baselines
In all datasets, we evaluate the performance of RTMPose (Jiang et al., 2023), HRNet (Sun et al., 2019), and SwinPose (Xiong et al., 2022) to estimate the human pose. These models are common baseline models for top-down 2D pose estimation with diverse architectures.
@@
We use The Percentage of Detected Joints (PDJ) (Toshev and Szegedy, 2014), widely used for evaluating the accuracy of 2D human pose estimation. It measures the proportion of correctly detected keypoints based on a distance threshold relative to the torso diameter. This metric is useful for assessing pose estimation models with varying degrees of localization precision (Yeung et al., 2024). In this evaluation metric, a keypoint is deemed correctly detected if the normalized distance between the predicted and ground truth keypoints is less than 0.5. Furthermore, PDJ allows for the evaluation of pose estimation accuracy under different levels of localization precision by adjusting the threshold. The area under the PDJ curve (AUC) is computed to assess performance across various threshold values.
@@ 5. Experiments
5.1. Performance in Track-ID task
Experiments were conducted on the Track-ID task on fixed camera datasets, Indoor and Outdoor, as described above. We experimented with this task mainly on the in-play portion: Indoor was all frames of each video, while Outdoor excluded the out-of-play portion and divided the video of one game into several segments. Two types of video segments exist: one starting with a check-ball at the top and the other starting with a free throw (the former number of videos is 74, and the latter number of videos is 22). The numbers of video frames in Indoor and Outdoor datasets 179.38 $\pm$ 64.77 and 1202.84 $\pm$ 1025.92 (mean $\pm$ standard deviation), respectively. The evaluation metrics are TI-HOTA and its sub-scores TI-DetA, TI-AssA as described above. For the distance tolerance parameter, $\tau$ of TI-HOTA, $\tau=1$ was adopted given the ratio to the football pitch.
Tables 3 show the results, including the mean and standard deviation, when ID Switch Merging was performed and when it was not performed, demonstrating the effectiveness and robustness of the corresponding module. Note that ID Switch based on two identification information is not reported because it did not occur in the proposed baseline .
It is difficult to compare the Indoor and Outdoor results simply , but the better performance of Indoor can be explained to some extent by the shorter video length and the easier identification of attributes. This is because if ID exchange between players occurs in tracking (see Figure 3), the attributes will no longer be correct from then on, and the match between the prediction and the ground truth will fail for a large part of the video clip.
Figure 3. Exmaple of ID exchange during tracking in Outdoor dataset. It occurs during occlusion and is a significant contributing factor to the decrease in the TI‑HOTA score.
@@ 5.1. Performance in Track-ID task
Indoor RTMPose (Jiang et al., 2023) 68.6968.69% 87.2687.26% 71.9971.99% 70.0270.02% 77.8777.87% 78.2978.29% 75.6975.69% 37.8437.84%
HRNet (Sun et al., 2019) 69.5169.51% 86.7686.76% 68.6268.62% 66.3666.36% 77.3177.31% 75.6075.60% 74.0374.03% 37.0137.01%
SwinPose(Xiong et al., 2022) 69.2069.20% 86.1386.13% 64.1964.19% 67.4867.48% 77.2777.27% 74.4474.44% 73.1273.12% 36.5636.56%
Outdoor RTMPose (Jiang et al., 2023) 88.8288.82% 93.5793.57% 91.0391.03% 85.8685.86% 89.4089.40% 87.6087.60% 89.4389.43% 45.1245.12%
HRNet (Sun et al., 2019) 88.3988.39% 93.8893.88% 89.5389.53% 85.6585.65% 91.4391.43% 89.0989.09% 88.5188.51% 45.2145.21%
SwinPose(Xiong et al., 2022) 88.2188.21% 93.1493.14% 89.1989.19% 85.2085.20% 90.1690.16% 90.1690.16% 89.2789.27% 45.0345.03%
Drone RTMPose (Jiang et al., 2023) 85.5785.57% 85.6085.60% 85.1785.17% 80.6480.64% 77.5977.59% 84.3084.30% 83.4583.45% 42.0742.07%
HRNet (Sun et al., 2019) 85.9985.99% 86.0986.09% 84.5384.53% 79.6979.69% 84.6084.60% 84.7584.75% 84.0784.07% 42.3842.38%
SwinPose(Xiong et al., 2022) 85.4985.49% 86.3086.30% 84.4384.43% 79.7279.72% 83.0583.05% 84.7084.70% 83.8883.88% 42.2842.28%
@@
Table 4. 2D pose estimation models performance on all datasets, ranked by PDJ. The best-performing result for each dataset is highlighted in bold. Sho, El, Wri, Cen, and Ank represent the average PDJ of Shoulder, Elbow, Wrist, Center, and Ankle, respectively. The Center is defined as the midpoint between the left and right hips. The PDJ column represents the mean PDJ across all 10 keypoints. AUC represents the area under the PDJ curve as the threshold varies from 0 to 0.5 in increments of 0.01, with a maximum value of 0.5.
5.2. 2D Pose Estimation Performance
@@
We also evaluated the performance of 2D pose estimation models in our all datasets. We evaluated three models without fine-tuning as baselines. Since each player in our datasets is annotated with only 10 keypoints, while the baseline models output 17 keypoints in COCO (Lin et al., 2015) keypoints format, we select the corresponding 10 keypoints for evaluation to ensure consistency.
@@
The evaluation results are summarized in Table 4. All models achieved their highest PDJ scores on the Outdoor dataset, but performance declined on the Drone dataset, with the Indoor dataset exhibiting the lowest accuracy. The lower performance observed in the Indoor dataset can be primarily attributed to its lower image resolution (1280 × 720) compared to the Outdoor and Drone datasets (3820 × 2160), which may result in blurred keypoints and reduced model accuracy.
@@
A comparison between different keypoints across datasets, ankles, shoulders, and the center exhibit the most stable and high detection accuracy in all scenarios. In contrast, the elbows and wrists in the upper limbs demonstrate weaker performance, probably due to occlusion and motion blur. In our TrackID3x3 dataset, RTMPose(Jiang et al., 2023) achieved the best performance in the Outdoor dataset, with an average PDJ of 89.43%. This result is highly comparable to the average PDJ of 89.51% reported for RTMPose in AutoSoccerPose (Yeung et al., 2024). For illustration, Figure 4 presents examples of players with lower prediction accuracy in the three datasets.
@@
Figure 4. Examples of players with low PDJ across three datasets. From left to right, the images correspond to the Indoor, Outdoor, and Drone datasets. The players positioned at the back are the primary targets for keypoint detection. However, due to occlusion, keypoint prediction for the obstructed player results in partial or complete failure.
6. Conclusion
@@
Although we were unable to provide quantitative or qualitative evidence, tracking methods based on Tracking by Detection, such as CAMELTrack, may actually increase mis-tracking during occlusion when using detectors with excessively high detection power. This is because the bounding boxes of hidden players may be nearly identical to those of players in the foreground (i.e., there are many pairs with very high IoU), which can easily lead to incorrect associations. This is particularly pronounced in sports like basketball, where occlusions occur frequently, and could be a contributing factor to ID exchange, one of the most critical challenges in the Track-ID task. The proposed baseline takes this trade-off into account and deliberately uses a detector with reduced detection power (though still sufficiently effective for tracking). However, ID exchange still occurred, so to truly address this issue, it may be necessary to adopt tracking algorithms that include ID Switch countermeasures during occlusion (such as OccluTrack (Gao et al., 2023)) or introduce post-processing methods that include tracklet splitting (such as GTA (Sun et al., 2024)). The latter currently uses only appearance features for split, so there may be cases where the split does not work properly when ID exchange occurs between teammates (because teammates often have similar appearances) but by utilizing prediction results of jersey number, it may be possible to achieve more effective splitting and merging, potentially bringing us closer to resolving this issue.
Acknowledgements. This work was financially supported by JST PRESTO Grant Number JPMJPR20CA, JSPS KAKENHI Grant Number 23H03282, JP24K23889 and Mizuno Sports Promotion Foundation.
Appendix A MOT experiment on Drone dataset
We describe a simple baseline for player tracking in Drone dataset. We implement two conventional baseline MOT algorithms: ByteTrack (Zhang et al., 2022) and BoT-SORT (Aharon et al., 2022).
For performance evaluation, we use HOTA (Luiten et al., 2021) metric due to its comprehensive ability to measure both detection and association performance. Additionally, we incorporate supplementary metrics, including DetA, AssA, and ID switches (IDs), to provide a more granular analysis of tracking performance.
Table A.4 shows the results with the default parameters of the two methods. The results reveal that the BoT-SORT method demonstrates improvement in HOTA scores compared to ByteTrack, along with a reduction in ID switches. However, there remains considerable room for further performance enhancement due to many occlusions.
Method Length [s] HOTA(%)↑ DetA(%)↑ AssA(%)↑ IDs↓
ByteTrack (Zhang et al., 2022) 39.55 $\pm$ 36.30 42.92 $\pm$ 11.96 37.02 $\pm$ 8.86 51.77 $\pm$ 22.21 15.06 $\pm$ 14.22
BoT-SORT (Aharon et al., 2022) 39.55 $\pm$ 36.30 49.98 $\pm$ 12.24 42.22 $\pm$ 9.63 61.82 $\pm$ 22.46 11.81 $\pm$ 13.19
Table A.4. ByteTrack and BoT-SORT results on Drone dataset.